# arXiv:2601.01248v3[math.OC] 02 Mar 2026

### v3

@@ 1.1 Optimization over finite-dimensional spaces: $\mathcal{X}=\mathbb{R}^{d}$
Here and below, the $\mathbb{R}^{d}$-valued initial value $x_{0}$ and the $d$-dimensional Wiener process $(W_{t})_{t\geq 0}$ are defined on a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ so that the filtration $(\mathcal{F}_{t})_{t\geq 0}$ is generated by $x_{0}$ and $(W_{t})_{t\geq 0}$ and augmented by all the $\mathbb{P}$-null sets in $\mathcal{F}$. In particular, $x_{0}$ is $\mathcal{F}_{0}$-measurable and independent of the Wiener process $(W_{t})_{t\geq 0}$. The admissible control $\theta$ is an $(\mathcal{F}_{t})_{t\geq 0}$-adapted process, and the set of admissible controls, denoted by $\Theta$, consists of all such control processes on $[0,1]$ that are square-integrable on $\Omega\times[0,1]$
@@ 1.2.2 Main result on convergence analysis as $\varepsilon\to 0$ and $N\to\infty$
$\left|\frac{1}{N}v^{N}_{\varepsilon}(0,\mathbf{x}^{0})-\mathcal{V}_{0}\right|\leq\underbrace{\frac{L_{\varepsilon}}{2N}}_{\text{Particle Error}}+\underbrace{C_{2}\cdot\varepsilon\ln\left(\frac{1}{\varepsilon}\right)}_{\text{Regularization Error}},$ (1.26)
where $\mathbf{x}^{0}=(x^{0}_{1},\ldots,x^{0}_{N})=(x_{0},\ldots,x_{0})\in\mathbb{R}^{Nd}$, $L_{\varepsilon}$ depends on $\varepsilon$ and the regularity of $G$, and $C_{2}$ depends on $d$, the Lipschitz constant of $G$, and the 2-Wasserstein distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between $\delta_{x_{0}}$ and $\mu^{*}$.
The above theorem summarizes the convergence results presented in Theorems 3.3 and 3.5. For clarity of exposition, we assume that all particles start from the same deterministic point $x_{0}\in\mathbb{R}^{d}$. More generally, if the initial positions are sampled i.i.d. according to an arbitrary law $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then the total approximation error contains an additional term that accounts for the discrepancy between $\mu_{0}$ and the empirical initial distribution; see Corollary 3.4. Moreover, under the regularity, convexity, and coercivity conditions of Proposition 3.2, the constant $L_{\varepsilon}$ in the particle approximation estimate can be chosen uniformly in $\varepsilon$; see Theorem 3.3.
@@ 1.3 Literature, contributions, and organization of the paper
  • The optimization over probability measures is approximated through mean-field control problems which are further approximated via $N$-particle systems (equivalently, $N$-player potential games). We establish rigorous convergence results, demonstrating that the value of the controlled finite-particle system converges to the global minimum of the objective functional at the rate of $\frac{C_{\varepsilon}}{N}+\varepsilon\ln\left(\frac{1}{\varepsilon}\right)$ as the regularization parameter $\varepsilon$ approaches zero and the particle number $N$ tends to infinity.
@@ 2.3 Numerical approximations with examples
9:     Compute weights $w_{j}\leftarrow\exp\left(-\frac{G(X_{t}^{(i)}+\sqrt{\tau}Z_{j})}{\varepsilon}\right)$
@@ 3.1 Finite Particle Approximation
    (Convexity): For each $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the total Hessian operator $\mathcal{H}_{G}(\mu)$, defined by its action on perturbations $\phi\in L^{2}(\mathbb{R}^{d},\mu;\mathbb{R}^{d})$ via the quadratic form
    $\langle\mathcal{H}_{G}(\mu)\phi,\phi\rangle:=\mathbb{E}\left[\langle\partial_{x}\partial_{\mu}G(\mu)(X)\cdot\phi(X),\phi(X)\rangle\right]+{\mathbb{E}}\left[\langle\partial^{2}_{\mu\mu}G(\mu)(X)(\tilde{X})\cdot\phi(\tilde{X}),\phi(X)\rangle\right],$
    with $X,\tilde{X}$ being independent random variables with law $\mu$, is a positive semi-definite operator.
@@
Then the value function $\mathcal{V}_{\varepsilon}$ of the mean-field control problem (1.19) is the unique classical solution to the master equation (1.20) (as in Assumption 3.1). Moreover, $\mathcal{V}_{\varepsilon}$ satisfies the following $\varepsilon$-uniform estimates:
$|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)|\leq\sqrt{d}K_{G},\quad|\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)(y)|\leq\frac{(K_{G})^{2}}{\kappa}+K_{G},$ (3.1)
for all $(t,\mu,x,y)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d}$.
@@ 3.3 Numerical approximation with examples
The pseudocode is presented in Algorithm 2 which we refer to as the Stochastic Control Method (SCM) for Optimization on $\mathcal{P}_{2}(\mathbb{R}^{d})$. The output is an empirical measure approximating the optimal distribution.
@@
6:   Project Context: $\tilde{X}^{(k)}\leftarrow X_{t}^{(k)}+\sqrt{\tau}\zeta^{(k)}$ for $\zeta^{(k)}\sim\mathcal{N}(0,I)$; form $\tilde{\mu}_{T}\approx\frac{1}{N}\sum\delta_{\tilde{X}^{(k)}}$
7:   for $i=1$ to NN do
8:     Sample $Y_{i,j}\leftarrow X_{t}^{(i)}+\sqrt{\tau}\xi_{i,j}$, for $j=1..S$
9:     Compute weights $w_{i,j}\propto\exp\left(-NG(Y_{i,j};\tilde{\mu}_{T})/\varepsilon\right)$
10:     Compute drift $\hat{\theta}_{i}\leftarrow\frac{1}{\tau}\left(\frac{\sum w_{i,j}Y_{i,j}}{\sum w_{i,j}}-X_{t}^{(i)}\right)$
11:     Update $X_{t+\Delta t}^{(i)}\leftarrow X_{t}^{(i)}+\hat{\theta}_{i}\Delta t+\sqrt{\Delta t}Z^{(i)}$, for $Z^{(i)}\sim\mathcal{N}(0,I)$
12:   end for
13:  end for
14:  $X_{0}^{(i)}\leftarrow X_{T}^{(i)}$ $\triangleright$ Recycle terminal state
15:end for
16:Output: $\mu^{*}\approx\frac{1}{N}\sum\delta_{X_{T}^{(i)}}$
In the computation of $\beta^{i}$ in (3.11) or the weights in Algorithm 2, the evaluation of the functional $G$ on empirical measures is required. This can be done efficiently for various classes of functionals, such as integral functionals and interaction energies. For example, if
$G(\mu)=\int_{\mathbb{R}^{d}}F(x)\,d\mu(x)+\frac{1}{2}\iint_{\mathbb{R}^{2d}}\Phi(x-y)\,d\mu(x)\,d\mu(y),$ (3.12)
then for an empirical measure $\mu^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}$, we have
$G(\mu^{N})=\frac{1}{N}\sum_{i=1}^{N}F(x_{i})+\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\Phi(x_{i}-x_{j}).$ (3.13)
Then for each particle $i$, the independence of noises gives us the following simplified computation
$\beta^{i}(t_{k},X_{t_{k}}^{i})\approx\frac{\sum_{l=1}^{S}\exp\left(-\frac{1}{\varepsilon}\left(F(Y_{i,l})+\frac{1}{2N}\sum_{j=1}^{N}\Phi(Y_{i,l}-Y_{j,l})+\frac{1}{2N}\sum_{j\neq i}\Phi(Y_{j,l}-Y_{i,l})\right)\right)Y_{i,l}}{\sum_{l=1}^{S}\exp\left(-\frac{1}{\varepsilon}\left(F(Y_{i,l})+\frac{1}{2N}\sum_{j=1}^{N}\Phi(Y_{i,l}-Y_{j,l})+\frac{1}{2N}\sum_{j\neq i}\Phi(Y_{j,l}-Y_{i,l})\right)\right)},$ (3.14)
where $Y_{j,l}=\sqrt{1-t_{k}}\xi_{j,l}+X_{t_{k}}^{j}$, for $j=1,\dots,N$ and $l=1,\dots,S$.
@@
In the above 2D Newtonian swarm example, it is not appropriate to study the sensitivity with respect to the regularization parameter $\varepsilon$ because the particle approximation error also depends on $\varepsilon$; see Theorem 3.3. To isolate the regularization error, we consider the following example that satisfies the regularity, convexity, and coercivity assumptions in Proposition 3.2 under which the constant $L_{\varepsilon}$ in Thoerem 3.3 is independent of $\varepsilon$.
@@
$V(x)$ $=\min\left(\frac{1}{2}\big(\|x-c_{L}\|^{2}-R^{2}\big)^{2},\quad\frac{1}{2}\big(\|x-c_{R}\|^{2}-R^{2}\big)^{2}\right)\,$,

### v4

# Stochastic Control Methods for Optimization

@@ 1.1 Optimization over finite-dimensional spaces: $\mathcal{X}=\mathbb{R}^{d}$
Here and below, we take terminal time $T=1$ and the $\mathbb{R}^{d}$-valued initial value $x_{0}$ and the $d$-dimensional Wiener process $(W_{t})_{t\geq 0}$ are defined on a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbb{P})$ so that the filtration $(\mathcal{F}_{t})_{t\geq 0}$ is generated by $x_{0}$ and $(W_{t})_{t\geq 0}$ and augmented by all the $\mathbb{P}$-null sets in $\mathcal{F}$. In particular, $x_{0}$ is $\mathcal{F}_{0}$-measurable and independent of the Wiener process $(W_{t})_{t\geq 0}$. The admissible control $\theta$ is an $(\mathcal{F}_{t})_{t\geq 0}$-adapted process, and the set of admissible controls, denoted by $\Theta$, consists of all such admissible control processes $\theta$ satisfying $\mathbb{E}\left[\left(\int_{0}^{1}|\theta_{s}|ds\right)^{2}\right]<\infty$.
@@ 1.2.2 Main result on convergence analysis as $\varepsilon\to 0$ and $N\to\infty$
$\left|\frac{1}{N}v^{N}_{\varepsilon}(0,\mathbf{x}^{0})-\mathcal{V}_{0}\right|\leq\underbrace{\frac{L_{\varepsilon}}{2N}}_{\text{Particle Error}}+\underbrace{C\cdot\varepsilon\ln\left(\frac{1}{\varepsilon}\right)}_{\text{Regularization Error}},$ (1.26)
where $\mathbf{x}^{0}=(x^{0}_{1},\ldots,x^{0}_{N})=(x_{0},\ldots,x_{0})\in\mathbb{R}^{Nd}$, $L_{\varepsilon}$ depends on $\varepsilon$ and the regularity of $G$, and the constant $C$ depends on $d$, the Hölder norm and exponent of $G$, and the 2-Wasserstein distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between $\delta_{x_{0}}$ and $\mu^{*}$.
The above theorem summarizes the convergence results presented in Theorems 3.3 and 3.5. For clarity of exposition, we assume that all particles start from the same deterministic point $x_{0}\in\mathbb{R}^{d}$. More generally, if the initial positions are sampled i.i.d. according to an arbitrary law $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then the total approximation error contains an additional term that accounts for the discrepancy between $\mu_{0}$ and the empirical initial distribution; see Corollary 3.4. Moreover, under the regularity, averaged convexity, and coercivity conditions of Proposition 3.2, the constant $L_{\varepsilon}$ in the particle approximation estimate can be chosen uniformly in $\varepsilon$; see Theorem 3.3.
@@ 1.3 Literature, contributions, and organization of the paper
  • The optimization over probability measures is approximated through mean-field control problems which are further approximated via $N$-particle systems (equivalently, $N$-player potential games). We establish rigorous convergence results, demonstrating that the value of the controlled finite-particle system converges to the global minimum of the objective functional at the rate of $\frac{L_{\varepsilon}}{N}+C\varepsilon\ln\left(\frac{1}{\varepsilon}\right)$ as the regularization parameter $\varepsilon$ approaches zero and the particle number $N$ tends to infinity.
@@ 2 Optimization over finite-dimensional spaces: $\mathcal{X}=\mathbb{R}^{d}$
For a matrix $A$, we use $|A|$ to denote its Frobenius/Hilbert-Schmidt norm, while by $|A|_{op}$ we denote its operator norm that is its largest singular value.
@@ 2.3 Numerical approximations with examples
9:     Compute weights $w_{j}\leftarrow\exp\left(-\frac{G(X_{t}^{(i)}+\sqrt{\tau}Z_{j})-\min_{k}G(X_{t}^{(i)}+\sqrt{\tau}Z_{k})}{\varepsilon}\right)$
@@ 3.1 Finite Particle Approximation
    (Averaged convexity): For each $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and every pair $(X,\eta)$ of $\mathbb{R}^{d}$-valued square integrable random variables with $\mathcal{L}(X)=\mu$,
    $\mathbb{E}\left[\frac{1_{|\eta|>0}}{|\eta|}\langle\eta,\partial_{x}\partial_{\mu}G(\mu)(X)\eta\rangle+\frac{1_{|\eta|>0}}{|\eta|}\Big\langle\eta,\widetilde{\mathbb{E}}\big[\partial^{2}_{\mu\mu}G(\mu)(X)(\widetilde{X})\widetilde{\eta}\big]\Big\rangle\right]\geq 0$
[…]
@@
Then the value function $\mathcal{V}_{\varepsilon}$ of the mean-field control problem (1.19) is the unique classical solution to the master equation (1.20) (as in Assumption 3.1). Moreover, $\mathcal{V}_{\varepsilon}$ satisfies the following $\varepsilon$-uniform estimates: for all $(t,\mu,x)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$
$|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)|_{op}\leq K_{G},$
$|\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)(y)|_{op}\leq\frac{(K_{G})^{2}}{\kappa}+K_{G},\quad\mu\text{-a.e. }y\in\mathbb{R}^{d}.$ (3.1)
@@ 3.3 Numerical approximation with examples
The main computational difficulty is that the global cloud weight in (3.11) depends on the full random terminal configuration $\frac{1}{N}\sum_{j=1}^{N}\delta_{X^{j}_{t_{k}}+\sqrt{1-t_{k}}\xi^{j}}$. Thus, at every time step and for every Monte Carlo sample, one must evaluate $G$ on an $N$-particle empirical measure, and the resulting Gibbs factor $\exp\!\left(-\frac{N}{\varepsilon}G\!\left(\frac{1}{N}\sum_{j=1}^{N}\delta_{X^{j}_{t_{k}}+\sqrt{1-t_{k}}\xi^{j}}\right)\right)$ accumulates fluctuations from all particles simultaneously. When $N$ is large or $\varepsilon$ is small, this produces severe weight concentration and a substantial increase in computational cost, especially for interacting functionals $G$ whose evaluation already scales superlinearly in $N$.
For a general functional $G:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$, a natural particlewise localization is obtained by freezing the background empirical measure and varying only one particle. More precisely, for a background measure $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and a candidate position $y\in\mathbb{R}^{d}$, define the particlewise effective cost
$\mathcal{G}_{N}(y;\nu):=N\,G\!\left(\frac{1}{N}\delta_{y}+\frac{N-1}{N}\nu\right).$ (3.12)
Assume that $G$ admits a linear functional derivative $\frac{\delta G}{\delta\mu}(\nu)(\cdot)$ and a second-order expansion at $\nu$ in the direction of signed perturbations. Since $\frac{\delta G}{\delta\mu}(\nu)(\cdot)$ is defined only up to an additive constant, the combination appearing below is invariant under the choice of representative. Writing $\eta_{N}^{y,\nu}=\frac{1}{N}(\delta_{y}-\nu)$, so that $\frac{1}{N}\delta_{y}+\frac{N-1}{N}\nu=\nu+\eta_{N}^{y,\nu}$, we obtain formally
$\mathcal{G}_{N}(y;\nu)=NG(\nu)+\left[\frac{\delta G}{\delta\mu}(\nu)(y)-\int_{\mathbb{R}^{d}}\frac{\delta G}{\delta\mu}(\nu)(z)\,\nu(dz)\right]+O\!\left(\frac{1}{N}\right).$ (3.13)
Indeed,
$G(\nu+\eta_{N}^{y,\nu})=G(\nu)+\int_{\mathbb{R}^{d}}\frac{\delta G}{\delta\mu}(\nu)(z)\,\eta_{N}^{y,\nu}(dz)+O\!\left(\frac{1}{N^{2}}\right),$
and multiplying by $N$ yields (3.13). Hence the dominant term $NG(\nu)$ is independent of $y$ and cancels in the ratio defining $\beta^{i}$, while the remaining fluctuation is only of order $O(1)$. This suggests particle-wise approximation
$\beta^{i}(t_{k},X_{t_{k}}^{i})\approx\frac{\mathbb{E}\!\left[\exp\!\left(-\frac{1}{\varepsilon}\mathcal{G}_{N}\big(\sqrt{1-t_{k}}\,\xi^{i}+x^{i};\,\eta^{i}\big)\right)\,\big(\sqrt{1-t_{k}}\,\xi^{i}+x^{i}\big)\right]}{\mathbb{E}\!\left[\exp\!\left(-\frac{1}{\varepsilon}\mathcal{G}_{N}\big(\sqrt{1-t_{k}}\,\xi^{i}+x^{i};\,\eta^{i}\big)\right)\right]}\bigg|_{x^{i}=X_{t_{k}}^{i},\,\eta^{i}=\nu_{t_{k}}^{N,-i}},$ (3.14)
where $\nu_{t_{k}}^{N,-i}:=\frac{1}{N-1}\sum_{j\neq i}\delta_{\widehat{Y}^{(i)}_{j}}$ with $\widehat{Y}^{(i)}_{j}$ being a sample of $X^{j}_{t_{k}}+\sqrt{1-t_{k}}\,\xi^{j}$ at time $t_{k}$ associated to particle $i$, is the empirical background seen by particle $i$. Heuristically, this is a frozen-background or propagation-of-chaos approximation with the other particles being treated through their empirical law. Note that even when $G$ is not smooth enough for the first-order expansion (3.13), $\mathcal{G}_{N}(y;\nu)$ remains a natural expression for isolating the local contribution of one particle and the above particle-wise approximation works.
In practice, one often employs a broader localized family of approximations. First, the background measure $\nu_{t_{k}}^{N,-i}$ may be replaced by a frozen proxy $\widehat{\nu}_{t_{k}}^{N,-i}$, for instance the current empirical swarm, a noisy forward projection of that swarm, or another cheaply computed proxy of the future crowd. Second, when evaluating $\mathcal{G}_{N}(y;\widehat{\nu}_{t_{k}}^{N,-i})$ remains expensive, one may replace the full background measure by a subsampled empirical one, i.e.,
$\widehat{\nu}_{t_{k}}^{N,-i,m}:=\frac{1}{m}\sum_{r=1}^{m}\delta_{\widehat{Y}_{J_{r}}^{(i)}},$ (3.15)
where $J_{1},\dots,J_{m}$ are sampled indices from the global cloud and $\widehat{Y}_{J_{r}}^{(i)}$ denotes a sample of $X^{J_{r}}_{t_{k}}+\sqrt{1-t_{k}}\,\xi^{J_{r}}$ at time $t_{k}$, associated to particle $i$. Then we use the quantity
$\mathcal{G}_{N,m}(y;\widehat{\nu}_{t_{k}}^{N,-i}):=N\,G\!\left(\frac{1}{N}\delta_{y}+\frac{N-1}{N}\widehat{\nu}_{t_{k}}^{N,-i,m}\right),$ (3.16)
or, more generally, any computational proxy $\widetilde{\mathcal{G}}_{N,m}(y;\widehat{\nu}_{t_{k}}^{N,-i})$ to approximate the local contribution of particle $i$. The corresponding Monte Carlo update then takes the generic form
$\beta^{i}(t_{k},X_{t_{k}}^{i})\approx\frac{\sum_{l=1}^{S}\exp\!\left(-\frac{1}{\varepsilon}\widetilde{\mathcal{G}}_{N,m}(Y_{i,l};\widehat{\nu}_{t_{k}}^{N,-i})\right)Y_{i,l}}{\sum_{l=1}^{S}\exp\!\left(-\frac{1}{\varepsilon}\widetilde{\mathcal{G}}_{N,m}(Y_{i,l};\widehat{\nu}_{t_{k}}^{N,-i})\right)},$ (3.17)
where $Y_{i,l}=X_{t_{k}}^{i}+\sqrt{1-t_{k}}\,\xi_{i,l}$. The case m=N1m=N-1 and $\widetilde{\mathcal{G}}_{N,m}=\mathcal{G}_{N}$ recovers the global cloud weight with a frozen background measure, while smaller $m$ yields a further variance-cost tradeoff through subsampling. The four numerical examples considered below are all of this localized type. They all replace the global cloud weight by a particle-wise one, and the global cloud is subsampled when evaluating the interaction term, conditionally in Example 3.7 and explicitly in Examples 3.6, 3.8 and 3.9.
This type of localization is consistent with several established ideas in the literature. In high-dimensional sequential Monte Carlo, local or block particle filters replace global weights by localized weights precisely to avoid weight collapse and the curse of dimensionality; see [56, 61]. From a statistical viewpoint, the same structural principle underlies mean-field variational approximations and particle/coordinate-wise updates for Gibbs-type models; see [38]. In our setting, the particle-wise approximation and its subsampled variants play the analogous role of replacing the full interacting cloud weight by a one-particle effective weight against a frozen background measure.
For clarity, Algorithm 2 below is presented in its exact global cloud weight form, which discretizes as in (3.11). This exact version serves as the conceptual reference scheme. In actual implementations, one may replace the global weight in (3.11) by the particle-wise approximation (3.17) whenever the exact evaluation of $G$ becomes prohibitively expensive or numerically unstable. The pseudocode is presented in Algorithm 2 which we refer to as the Stochastic Control Method (SCM) for Optimization on $\mathcal{P}_{2}(\mathbb{R}^{d})$. The output is an empirical measure approximating the optimal distribution.
@@
6:   for $j=1$ to $S$ do
7:     Sample $\xi_{1,j},\dots,\xi_{N,j}\sim\mathcal{N}(0,I_{d})$ independently
8:     Set $Y_{r,j}\leftarrow X_{t}^{(r)}+\sqrt{\tau}\,\xi_{r,j}$, for $r=1,\dots,N$
9:     Form $\mu_{T}^{(j)}\leftarrow\frac{1}{N}\sum_{r=1}^{N}\delta_{Y_{r,j}}$
10:     Compute weight $w_{j}\leftarrow\exp\left(-\frac{N}{\varepsilon}G\left(\mu_{T}^{(j)}\right)\right)$
11:   end for
12:   for $i=1$ to NN do
13:     Compute $\beta_{i}\leftarrow\frac{\sum_{j=1}^{S}w_{j}Y_{i,j}}{\sum_{j=1}^{S}w_{j}}$
14:     Compute drift $\hat{\theta}_{i}\leftarrow\frac{\beta_{i}-X_{t}^{(i)}}{\tau}$
15:     Update $X_{t+\Delta t}^{(i)}\leftarrow X_{t}^{(i)}+\hat{\theta}_{i}\Delta t+\sqrt{\Delta t}Z^{(i)}$, for $Z^{(i)}\sim\mathcal{N}(0,I)$
16:   end for
17:  end for
18:  $X_{0}^{(i)}\leftarrow X_{T}^{(i)}$ $\triangleright$ Recycle terminal state
19:end for
20:Output: $\mu^{*}\approx\frac{1}{N}\sum\delta_{X_{T}^{(i)}}$
To bridge the gap between the idealized formulation in Algorithm 2 and its numerical implementations, some other techniques are incorporated. Although the pseudocode assumes a fixed time step $\Delta t=T/M$ and a constant number of iterations $L$, practical applications may adopt adaptive, multi-phase scheduling, beginning with a standard time horizon ($T=1$) for global exploration and followed by refinement phases over shorter horizons (e.g., T=0.001) with finer time steps (e.g., $\Delta t=10^{-6}$), thereby enhancing stability during convergence to complex target distributions. Furthermore, direct evaluation of exponential weights $\exp(-\mathcal{C}/\varepsilon)$ is numerically infeasible for extremely small $\varepsilon$ (e.g., $10^{-299}$) due to floating-point underflow; this is mitigated by employing the Log-Sum-Exp technique as in Algorithm 1, which performs computations in the logarithmic domain and normalizes exponents prior to exponentiation, ensuring numerical stability in near-zero regularization regimes.
@@
In the above 2D Newtonian swarm example, it is not appropriate to study the sensitivity with respect to the regularization parameter $\varepsilon$ because the particle approximation error also depends on $\varepsilon$; see Theorem 3.3. To isolate the regularization error, we consider the following example that satisfies the regularity, averaged convexity, and coercivity assumptions in Proposition 3.2 under which the constant $L_{\varepsilon}$ in Thoerem 3.3 is independent of $\varepsilon$.
@@
$V(x)$ $=\min\left(\frac{1}{2}\big(|x-c_{L}|^{2}-R^{2}\big)^{2},\quad\frac{1}{2}\big(|x-c_{R}|^{2}-R^{2}\big)^{2}\right)\,$,